A 55000.kg truck rounds a 38.0 m radius curve at 62.0km/h. Find how much centripetal force must be supplied by friction with the road to keep the truck from from slipping

1
Expert's answer
2022-02-28T10:38:13-0500

The centripetal force

"F=\frac{mv^2}{r}\\
=\rm\frac{55000\: kg*(17.2\: m/s)^2}{38.0\: m}=4.29*10^5\: N"